# A Roadmap to Pluralistic Alignment

### v1

@@ A.1 Further Analysis
As you can see in our results in Table 1, almost all pre-aligned models are more similar to the target human distribution than the post-aligned models for both datasets. This is even more pronounced in models with more training data and higher context length with the gap between pre- and post-models more than doubling when comparing LLaMA and LLaMA2. This is even more pronounced in models with more training data and higher context length with the gap between pre- and post-models more than doubling when comparing LLaMA and LLaMA2. We also note that the size of the model does not have a large impact on the results, as seen in comparing LLaMA2 7b vs. 13b. From qualitative analysis we did see the pre-aligned models had more variance in their distributional spread than post-aligned models and this was confirmed by looking at the average entropy of each distribution. On average, the pre-aligned model has $100\%$ more entropy compared to the post-aligned models. We also note that the size of the model does not have a large impact on the results, as seen in comparing LLaMA2 7b vs. 13b. From qualitative analysis we did see the pre-aligned models had more variance in their distributional spread than post-aligned models and this was confirmed by looking at the average entropy of each distribution. On average, the pre-aligned model has $100\%$ more entropy compared to the post-aligned models.
@@ Appendix B Additional Experimentation
In section 5.2 we explore the claim that pre-aligned models might perform better in superposition-pluralism than post-RLHF models. We test this hypothesis using two datasets, GlobalOpinnionQA and the Machine Personality Inventory. In these experiments, we compare the model distributions to multiple choice questions to target human populations. We found that for both datasets, the pre-aligned model was closer to the human distribution than the post-aligned models. From qualitative analysis we noticed that in the majority of cases the distributions for the pre-aligned models were more variable across the answer choices, in contrast to the post-aligned models which should more spiked distributions with probability mass centered on only one or two answer choices. This was reflected in our analysis of entropy, which showed that all pre-aligned models had higher average entropy across their distributions than post-aligned models. See Table 4 and Figure 3 for these results.
Model Human LLaMA LLaMA2 (7B) LLaMA2 (13B) GPT-3
Dataset Japan/US Global Pre Alpaca Tulu Pre Post Pre Post Pre Post
GlobalQA 0.96/0.99 NA 1.38 1.15 0.67 1.20 0.61 1.19 0.51 1.24 0.76
MPI NA 1.23 1.40 1.02 0.78 1.04 0.65 1.22 0.73 0.82 0.90
@@
Model Class LLaMA LLaMA2 (7B) LLaMA2 (13B) GPT-3
Dataset Pre Alpaca Tulu Pre Post Pre Post Pre Post
GlobalQA (Japan) 0.45 0.51 0.62 0.51 0.67 0.51 0.68 0.50 0.59
GlobalQA (US) 0.45 0.50 0.62 0.51 0.66 0.51 0.67 0.50 0.59
MPI 0.34 0.47 0.54 0.50 0.55 0.42 0.53 0.55 0.53

### v2

@@ A.1 Further Analysis
As you can see in our results in Table 1, almost all pre-aligned models are more similar to the target human distribution than the post-aligned models for both datasets. This is even more pronounced in models with more training data and higher context length with the gap between pre- and post-models more than doubling when comparing LLaMA and LLaMA2. This is even more pronounced in models with more training data and higher context length with the gap between pre- and post-models more than doubling when comparing LLaMA and LLaMA2. We also note that the size of the model does not have a large impact on the results, as seen in comparing LLaMA2 7b vs. 13b. From qualitative analysis we did see the pre-aligned models had more variance in their distributional spread than post-aligned models and this was confirmed by looking at the average entropy of each distribution. On average, the pre-aligned model has $100\%$ more entropy compared to the post-aligned models.
@@ Appendix B Additional Experimentation
In section 5.2 we explore the claim that pre-aligned models might perform better in distributional pluralism than post-RLHF models. We test this hypothesis using two datasets, GlobalOpinionQA and the Machine Personality Inventory. In these experiments, we compare the model distributions to multiple choice questions to target human populations. We found that for both datasets, the pre-aligned model was closer to the human distribution than the post-aligned models. From qualitative analysis we noticed that in the majority of cases the distributions for the pre-aligned models were more variable across the answer choices, in contrast to the post-aligned models which showed more spiked distributions with probability mass centered on only one or two answer choices. This was reflected in our analysis of entropy, which showed that all pre-aligned models had higher average entropy across their distributions than post-aligned models. See Table 4 and Figure 3 for these results.
Model Human LLaMA LLaMA2 (7B) LLaMA2 (13B) Gemma (7B) GPT-3
Dataset Japan/US/Germany Global Pre Alpaca Tulu Pre Post Pre Post Pre Post Pre Post
GlobalQA 0.96/0.99/0.96 NA 1.38 1.15 0.67 1.20 0.61 1.19 0.51 1.4 0.81 1.24 0.76
MPI NA 1.23 1.40 1.02 0.78 1.04 0.65 1.22 0.73 1.47 0.41 0.82 0.90
@@
Model Class LLaMA LLaMA2 (7B) LLaMA2 (13B) Gemma (7B) GPT-3
Dataset Pre Alpaca Tulu Pre Post Pre Post Pre Post Pre Post
GlobalQA (Japan) 0.45 0.51 0.62 0.51 0.67 0.51 0.68 0.45 0.61 0.50 0.59
GlobalQA (US) 0.45 0.50 0.62 0.51 0.66 0.51 0.67 0.45 0.61 0.50 0.59
GlobalQA (Germany) 0.46 0.57 0.63 0.52 0.69 0.53 0.68 0.47 0.61 0.51 0.62
MPI 0.34 0.47 0.54 0.50 0.55 0.42 0.53 0.33 0.68 0.55 0.53